Dexter bought a big crate of lanzones weighing 125 kg amounting to Php5,000.00. He repacked the lanzones and placed in a smaller crate with 2 1/2 kg per crate and sold each crate for Php135.00. How much was his gain?
Answer by Theo(13342)

the large crate of lanzones weighs 125 kg at a cost of 5,000 php.

he placed the lanzones into smaller crates weighing 2.5 kg each.

125 / 2.5 = 50 smaller crates.

50 * 135 php for each crate = 6750.

he paid 5000 and got 6750 back.

his profit is 6750 - 5000 = 1750

that was his gain.